‏Calculate the density , specific weight , and volume of chloride gas at 25 ° C and pressure of 600 000 N / m²abs


1
Expert's answer
2021-01-19T05:48:59-0500

Given the pressure of chloride gas at 25 ° C is

P=600000N/m2

and 1 pascal is equal to 1 N/m2.

therefore P= 600000 pa


Convert 25 ° C to k

T=273+25 =298K

Gas constant for chloride assume to be 117 kj/kg. K

take g to be g=9.81m/s2

(a) Density of chloride gas

From an ideal gas law P=ρRT

yo have P, R, T

density "\u03c1 =P\/RT"

ρ = 600000pa/(117 x103x298)

density=0.0172 kg/m3


(b) specific weight

weight=density x gravitational constant

w= ρ g

= 0.0172 X 9.81

w=0.1687 N/m3

(c) Volume

V=I/ ρ

V=1/0.0172

V=58.1395 m3/kg